OpenStudy (anonymous):

Giving your answer in interval notation, give the set of all x for which the expression (x+3x√x)/(x−6x√x) is defined.

OpenStudy (freckles):

\[\frac{x+3x \sqrt{x}}{x-6x \sqrt{x}}\] is that right?

OpenStudy (freckles):

You have a fraction so you need to make sure to exclude from your answer when the bottom is 0. You have square roots, we want to make sure the insides are positive or 0. Then w need to find the intersection of the sets we found from both of those sets I just mentioned.

OpenStudy (anonymous):

can you clarify? I'm confused by your explanation

OpenStudy (freckles):

find when \[x-6x*\sqrt{x}=0\] <--x cannot be these numbers also underneath the square roots should be positive or zero so we have x underneath the square root so we need x>=0